# GeMID: Generalizable Models for IoT Device Identification

### v1

@@ Abstract
With the proliferation of Internet of Things (IoT) devices, ensuring their security has become paramount. Device identification (DI), which distinguishes IoT devices based on their traffic patterns, plays a crucial role in both differentiating devices and identifying vulnerable ones, closing a serious security gap. However, existing approaches to DI that build machine learning models often overlook the challenge of model generalizability across diverse network environments. In this study, we propose a novel framework to address this limitation and evaluate the generalizability of DI models across datasets collected within different network environments. Our approach involves a two-step process: first, we develop a feature and model selection method that is more robust to generalization issues by using a genetic algorithm with external feedback and datasets from distinct environments to refine the selections. Second, the resulting DI models are then tested on further independent datasets in order to robustly assess their generalizability. We demonstrate the effectiveness of our method by empirically comparing it to alternatives, highlighting how fundamental limitations of commonly employed techniques such as sliding window and flow statistics limit their generalizability. Our findings advance research in IoT security and device identification, offering insights into improving model effectiveness and mitigating risks in IoT networks.
Index Terms:
IoT security, device identification, machine learning, generalizability.
I Introduction
@@
Despite the rapid proliferation of devices and providers, security remains a critical issue. IoT devices can be more challenging to secure than conventional devices due to their diverse hardware, software, and varied manufacturer profiles [2]. Moreover, unlike conventional devices, many IoT devices have non-standard interfaces which impair user interaction and make it difficult to mitigate against vulnerabilities [3]. The NETSCOUT Threat Intelligence Report [4] indicates that a new IoT device on the network typically faces its first attack within 5 hours and becomes a specific attack target within 24 hours, and that most attacks exploit vulnerabilities in IoT devices [5]. Beyond being targets in attacks, compromised IoT devices can serve as tools in botnet attacks, exemplified by Mirai, URSNIF, and BASHLITE [6], where captured devices are used to orchestrate high-volume Distributed Denial of Service (DDoS) attacks. While various methods have been developed to detect and mitigate ongoing attacks, primarily Intrusion Detection Systems (IDS), proactive measures such as device-specific updates, internet access restrictions or isolation can prevent these attacks before they happen. Given the variety, quantity and unfamiliar interfaces of IoT devices, it is impractical to rely on users to implement these solutions. As a result, DI systems have been developed to automatically identify devices based on their activity and make it possible to apply appropriate security measures to create a secure IoT ecosystem.
@@
In this study, we explicitly focus on developing DI model instances that are demonstrably generalizable across network environments. We measure this using multiple datasets containing the same devices operating in different network environments. We use a two-stage process to develop the models. The first stage involves feature and model selection, and focuses on identifying device-independent features that do not display over-dependence on their network environment. In the second stage, we use an independent dataset to train device-specific model instances. We then use a further dataset, containing the same devices operating in a different network environment, to evaluate the generalizability of these device-specific models.
@@
    Comprehensive evaluation of model generalizability: We provide an extensive assessment of ML-based DI models across diverse network environments, highlighting key factors that influence generalizability.
@@
    Novel study framework for assessing generalizability: We introduce a two stage framework specifically designed to evaluate the generalizability of DI models across varied conditions, offering a more rigorous methodology for testing model robustness.
@@
This paper is organized as follows: Section II reviews related work, Section III covers data selection and feature extraction, Section IV details the process used for selecting and evaluating features, Section V presents model selection and the results of our model generalizability study, Section VI discusses limitations, and Section VII concludes.
II Related Work
The field of IoT DI has been developing in earnest since 2017. Among the pioneering works is IoTSentinel [14], which builds models using 23 features extracted from packet headers. Packet headers, analogous to envelopes for data traveling over a network, contain information for routing and processing the data. By analyzing these headers, various useful features for network tasks can be extracted. Fig. 1 illustrates a sample network packet and lists the header features present in it. Extractable features from packet headers include source/destination IP/MAC addresses, protocol, Time-to-Live (TTL), flag information, port numbers, sequence/acknowledgment numbers, and checksum. A significant contribution of this study was the introduction of  Aalto University IoT device captures [15], one of the most widely used open-access datasets in the IoT DI domain. Subsequent studies, such as IoTSense [9], expanded the feature sets by incorporating payload features (such as payload size and entropy) alongside header-derived features. Concurrently, other research focused on statistics derived from headers, emphasizing the relationship between packets within a specified range (e.g., TTL [8], packet size [16], time [16]).
@@ II Related Work
The use of raw data is also possible. As shown in Fig. 1, a network packet can be represented as raw bytes. In this approach, the first $n$ bytes of the packet are provided as input to the ML models. Instead of being used explicitly as features, this data is fed into neural network-based learning models, in particular Convolutional Neural Networks (CNNs), where each byte is treated as an image pixel [21, 22, 23, 24, 25]. If the number of bytes exceeds $n$, truncation is applied; if less, padding is used. This method is conceptually similar to the packet header method; but its fields are not strictly separated.
@@
A more general issue with existing DI studies is the lack of cross-dataset validation, which involves evaluating a model trained on one dataset against a different dataset to assess its generalizability and robustness. While some studies use multiple datasets [11, 12, 13], they typically apply their methods separately to each, training and testing within the same dataset (e.g., training on a portion of Dataset A and testing on the remainder). This approach raises potential concerns about their performance in unseen environments. It also exacerbates the specific issues raised above surrounding the use of statistical features, since models using these features would — in effect — be tested within the same network environment as the one they were trained in. In this study, we demonstrate the importance of cross-dataset validation in developing generalizable models.
III Materials And Methods
III-A Data Selection
@@
In our study, we use data from two dataset families, UNSW and MonIoTr. Each family consists of multiple datasets that are useful for our model building and evaluation. We use the UNSW datasets for feature and model selection, and the MonIoTr datasets to test the robustness and generalizability of the resulting models.
III-A1 UNSW
@@
Notably, despite both datasets containing most devices in common (see Fig. 2), they were collected at different times, in distinct site environments with varying network characteristics, and for disparate purposes, likely by different users. We leverage these datasets for feature and model selection, employing UNSW-DI as training data and UNSW-AD as test data, and vice versa in different runs/iterations, to enhance model robustness and generalizability.
@@
III-A2 MonIoTr
After selecting the features and models, to remove any bias, we evaluate their effectiveness on a different dataset, IoT Information Exposure—IMC’19 (MonIoTr) [28]. The MonIoTr dataset consists of 81 devices. An important characteristic of this dataset is that data collection took place at two separate sites, one in the UK (MonIoTr-UK) and the other in the USA (MonIoTr-USA). Of the 81 devices, 26 devices were present in both sites, with the USA site having 47 devices (including 21 unique devices) and the UK site having 34 devices (including 8 unique devices). Fig. 3 shows the distribution of devices by country.
@@ III-A2 MonIoTr
Additionally, each device was connected via a VPN to the other site, enabling data collection across both sites for every condition. The same conditions were repeated using VPN tools to establish connections between the sites. For example, idle data was collected by connecting to the UK site from the USA via VPN and vice versa.
@@ III-B Feature Extraction
IV Feature Selection
During the feature selection phase, we aimed to identify the most effective features for DI. For this, we used the four UNSW data partitions described in Section III-A1 to guide feature selection. These comprise two sessions (S1 and S2) from each of UNSW-AD and UNSW-DI (referred to hereafter according to their dataset-session number: AD-S1, AD-S2, DI-S1, and DI-S2).
The generalizability of features is assessed in the following three ways, in order of increasing strictness: {LaTeXdescription}
@@ IV Feature Selection
IV-A Predictive Features
First, each feature is evaluated individually using each relevant combination of partitions, leading to 16 distinct evaluation contexts (see Fig. 4). Generalizability is assessed by training a decision tree (DT) model using the feature, with DT selected for speed and explainability. The utility of each feature is measured using the kappa metric, which adjusts for chance agreement.
@@ IV-A Predictive Features
Fig. 5 shows the resulting utility values. Notably, the scores measured under CV are much higher for some features than those measured under DD and SS scores. This indicates that information leaks may be occurring due to train and test folds being taken from the same partition, and suggests that CV is not a reliable basis for selecting features. Hence we only use DD and SS from now on.
@@
A voting system was used for feature selection, with each non-zero kappa value (with a tolerance of ±$\pm 5\%$, so $\geq 0.05$), indicating the feature had a positive contribution, resulting in a vote. Fig. 6 illustrates this system. Features receiving four or more votes from either DD (green) or SS (red) categories, with at least one DD vote, advanced to the next stage. Features not meeting this criterion were eliminated. The requirement for at least one DD vote is based on its greater strictness in assessing generalizability. As a result of voting, we identified 46 of the 332 features as being individually predictive.
@@
Figure 4: Data usage from UNSW datasets across different evaluation contexts. CV, SS, and DD steps are visualized in shades of blue, red, and green, respectively, throughout the paper for clarity. In the nomenclature, CV|AD-S1 denotes cross-validation on the AD dataset session 1. For cases other than CV, the first dataset is used for training and the second for testing. For example, AD-S1|DI-S2 means the first session of the UNSW-AD dataset is used for training and the second session of the UNSW-DI dataset is used for testing.
@@
Figure 5: Comparison of feature utility in UNSW datasets measured using CV (blue) and isolated methods, SS (red) and DD (green). CV tends to overestimate feature utility, with higher scores for many attributes. SS and DD produce more realistic evaluations. The discrepancy highlights the potential for information leakage in cross-validation and the importance of using isolated validation methods for assessing feature utility in ML-based DI models.
@@
Figure 6: Features identified as being predictive within one or more evaluation contexts (CV: blue, SS: red, DD: green), based on non-zero kappa scores of DT models. Note that the CV votes are not used to select features, but are shown here for completeness.
@@
Figure 7: List of intersecting features identified across eight dataset versus dataset (DD) cases using Genetic Algorithm (GA). Feature names highlighted in dark green were selected for model building after further post hoc analysis.
IV-B Feature Interactions

### v2

@@ Abstract
With the proliferation of devices on the Internet of Things (IoT), ensuring their security has become paramount. Device identification (DI), which distinguishes IoT devices based on their traffic patterns, plays a crucial role in both differentiating devices and identifying vulnerable ones, closing a serious security gap. However, existing approaches to DI that build machine learning models often overlook the challenge of model generalizability across diverse network environments. In this study, we propose a novel framework to address this limitation and to evaluate the generalizability of DI models across data sets collected within different network environments. Our approach involves a two-step process: first, we develop a feature and model selection method that is more robust to generalization issues by using a genetic algorithm with external feedback and datasets from distinct environments to refine the selections. Second, the resulting DI models are then tested on further independent datasets to robustly assess their generalizability. We demonstrate the effectiveness of our method by empirically comparing it to alternatives, highlighting how fundamental limitations of commonly employed techniques such as sliding window and flow statistics limit their generalizability. Moreover, we show that statistical methods, widely used in the literature, are unreliable for device identification due to their dependence on network-specific characteristics rather than device-intrinsic properties, challenging the validity of a significant portion of existing research. Our findings advance research in IoT security and device identification, offering insight into improving model effectiveness and mitigating risks in IoT networks.
keywords:
[…]
1 Introduction
@@
Despite the rapid proliferation of devices and providers, security remains a critical issue. IoT devices can be more challenging to secure than conventional devices due to their diverse hardware, software, and varied manufacturer profiles [2]. Moreover, unlike conventional devices, many IoT devices have nonstandard interfaces that impair user interaction and make it difficult to mitigate against vulnerabilities [3]. The NETSCOUT Threat Intelligence Report [4] indicates that a new IoT device on the network typically faces its first attack within 5 hours and becomes a specific attack target within 24 hours, and that most attacks exploit vulnerabilities in IoT devices [5]. Beyond being targets in attacks, compromised IoT devices can serve as tools in botnet attacks, exemplified by Mirai, URSNIF, and BASHLITE [6], where captured devices are used to orchestrate high-volume Distributed Denial of Service (DDoS) attacks. While various methods have been developed to detect and mitigate ongoing attacks, primarily Intrusion Detection Systems (IDS), proactive measures such as device-specific updates, internet access restrictions or isolation can prevent these attacks before they happen. Given the variety, quantity and unfamiliar interfaces of IoT devices, it is impractical to rely on users to implement these solutions. As a result, Device identification (DI) systems have been developed to automatically identify devices based on their activity and make it possible to apply appropriate security measures to create a secure IoT ecosystem. It is important to clarify that DI in this context aims to identify the device type (i.e., its make and model), and not to distinguish between unique instances of the same device type, which is a separate security challenge.
@@
A critical issue in current DI research is the reliance on flow and window statistics, which capture network environment dynamics rather than device-specific behavior. This study demonstrates that such statistical methods fail to generalize across diverse settings, undermining their reliability. In contrast, features from individual packets offer a robust alternative, as they are less influenced by external variables. This distinction has profound implications, potentially questioning the validity of numerous existing studies.
In this study, we explicitly focus on developing DI model instances that are demonstrably generalizable across network environments. We measure this using multiple datasets containing the same devices operating in different network environments. We use a two-stage process to develop the models. The first stage involves feature and model selection, and focuses on identifying device-independent features that do not display over-dependence on their network environment. Thus, while our work incorporates a sophisticated feature selection method, its purpose is to serve our primary goal: demonstrating a new, more robust methodology for creating and validating generalizable DI models.
In the second stage, we use an independent dataset to train device-specific model instances. We then use a further dataset, containing the same devices operating in a different network environment, to evaluate the generalizability of these device-specific models.
@@
    A novel framework for generalizable DI: We propose and validate a new framework for building and evaluating machine learning-based DI models that are generalizable across diverse network environments. This framework introduces a rigorous two-stage, cross-dataset validation methodology to ensure models are robust to real-world domain shifts.
@@
This paper is organized as follows: Section 2 reviews related work, Section 3 covers data selection and feature extraction, Section 4 details the process used for selecting and evaluating features, Section 5 presents model selection and the results of our model generalizability study, Section 6 discusses limitations, and Section 7 concludes.
2 Related Work
The field of IoT DI has been developing in earnest since 2017. Among the pioneering works is IoTSentinel [14], which builds models using 23 features extracted from packet headers. Packet headers, analogous to envelopes for data traveling over a network, contain information for routing and processing the data. By analyzing these headers, various useful features for network tasks can be extracted. Figure 1 illustrates a sample network packet and lists the header features present in it. Extractable features from packet headers include source/destination IP/MAC addresses, protocol, Time-to-Live (TTL), flag information, port numbers, sequence/acknowledgment numbers, and checksum. A significant contribution of this study was the introduction of  Aalto University IoT device captures [15], one of the most widely used open-access datasets in the IoT DI domain. Subsequent studies, such as IoTSense [9], expanded the feature sets by incorporating payload features (such as payload size and entropy) alongside header-derived features. Concurrently, other research focused on statistics derived from headers, emphasizing the relationship between packets within a specified range (e.g., TTL [8], packet size [16], time [16]).
@@ 2 Related Work
While flow-based features, such as flow volume and inter-arrival times, are prevalent [18, 19], they inherently encode network-specific traits—e.g., latency or bandwidth— rather than device characteristics. Studies like [17] and [20] exemplify this approach, yet their generalizability remains untested across diverse environments. Our analysis shows these methods are fragile, contrasting with packet-based features used in  [9, 14], which we validate as more reliable.
@@
Another approach moves away from hand-crafted features by using deep learning to automate the feature engineering process. A common technique is to use the raw bytes of a network packet directly, as illustrated in Figure 1. In this method, the first n bytes are fed into a model—often a Convolutional Neural Network (CNN)—where each byte is treated as an image pixel [21, 22, 23, 24, 25]. This data is truncated or padded to a fixed length n. This concept extends beyond raw packets to other data representations. For instance, DeviceMien [12] uses stacked LSTM-Autoencoders on TCP-flow data, while Lopez-Martin et al. [26] employ a hybrid CNN-RNN model to treat traffic flows as pseudo-images. Others focus on temporal patterns, such as using a CNN to learn long-term relationships from short-term packet arrival sequences [27].
@@
Beyond the methods that rely on flow statistics or concatenated packet headers, recent research has explored several other distinct avenues for IoT device identification.
One prominent approach focuses on the analysis of periodic background traffic, which is often characteristic of IoT device behavior. AUDI [29] and BehavIoT [30] exemplify this method. AUDI uses unsupervised learning, including Fourier transforms and signal autocorrelation, to model the periodic communication patterns of devices for identification. BehavIoT extends this by modeling not just periodic events but also user-initiated and aperiodic events to create comprehensive device and system behavior models from network traffic. These techniques capitalize on the insight that many IoT devices maintain constant, regular communication with backend servers, providing a stable basis for fingerprinting.
Another technique involves identifying fine-grained, packet-level signatures. The PINGPONG [31] system demonstrates that short, unique sequences of packet lengths and directions can serve as reliable fingerprints for specific device events (e.g., a light turning on). These signatures often appear as request-reply pairs (“ping-pong”) between a device and a cloud server and can be extracted automatically from traffic without deep statistical analysis.
Other methods have leveraged metadata from different network protocols. For instance, IoTFinder [32] uses passive DNS traffic, modeling a device’s identity based on the set of domain names it queries and the frequency of those queries. This DNS-based fingerprinting approach is designed to be scalable and effective even when devices are hidden behind Network Address Translation (NAT).
While these diverse techniques have proven effective within their respective evaluation contexts, a critical challenge that many early studies did not fully address is the generalizability of the resulting models. Work by Kolcun et al. [33] highlighted this issue by demonstrating that the accuracy of various ML models for device identification degrades significantly over time when tested on data collected weeks or months after the training period, suggesting that static models are insufficient. Complementing this, Ahmed et al. [34] conducted a comprehensive analysis across multiple datasets, showing that temporal, spatial, and data-collection-methodology differences all impact fingerprinting accuracy, and that features robust in one context may not be in another. These findings underscore the fundamental problem of model generalization, which is the primary focus of our work.
A related issue with existing DI studies is the lack of cross-dataset validation, which involves evaluating a model trained on one dataset against a different dataset to assess its generalizability and robustness. While some studies use multiple datasets [11, 12, 13], they typically apply their methods separately to each, training and testing within the same dataset (e.g., training on a portion of Dataset A and testing on the remainder). This approach raises potential concerns about their performance in unseen environments. It also exacerbates the specific issues raised above surrounding the use of statistical features, since models using these features would—in effect—be tested within the same network environment as the one they were trained in. In this study, we demonstrate the importance of cross-dataset validation in developing generalizable models.
Many of the studies cited in this section, such as [9, 10, 11, 12, 13, 14, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 28, 27, 31, 33, 30, 29, 32, 26, 12, 34] validate their approaches using training and testing sets drawn from the same dataset or environment. Our work diverges by focusing on the stricter challenge of model instance generalizability, where a specific model trained on data from one environment must perform accurately on data from a completely different environment. This evaluation distinction makes a direct F1-score comparison with these works an ‘apples-to-oranges’ comparison and is the reason our study implements its own baselines to ensure a fair and controlled experiment.
3 Materials And Methods
3.1 Data Selection
@@
In our study, we use data from two dataset families, UNSW and MonIoTr. Each family consists of multiple datasets that are useful for our model building and evaluation. We use the UNSW datasets for feature and model selection, and the MonIoTr datasets to test the robustness and generalizability of the resulting models. Additionally, we prioritized packet header features over flow or window statistics, as the latter are prone to network variability, reducing their utility for DI.
It is important to note that our framework uses these two dataset families for distinct stages (selection and evaluation) because they do not share a common set of device classes, as shown in Figures 2 and 3. Consequently, direct cross-testing, such as training a model on UNSW devices and testing on MonIoTr, is not possible as the classification labels are mutually exclusive.
3.1.1 UNSW
@@
Notably, despite both datasets containing most devices in common (see Figure 2), they were collected at different times, in distinct site environments with varying network characteristics, and for disparate purposes, likely by different users. We leverage these datasets for feature and model selection, employing UNSW-DI as training data and UNSW-AD as test data, and vice versa in different runs/iterations, to enhance model robustness and generalizability.
@@
3.1.2 MonIoTr
After selecting the features and models, to remove any bias, we evaluate their effectiveness on a different dataset, IoT Information Exposure—IMC’19 (MonIoTr) [36]. The MonIoTr dataset consists of 81 devices. An important characteristic of this dataset is that data collection took place at two separate sites, one in the UK (MonIoTr-UK) and the other in the USA (MonIoTr-USA). Of the 81 devices, 26 devices were present in both sites, with the USA site having 47 devices (including 21 unique devices) and the UK site having 34 devices (including 8 unique devices). Figure 3 shows the distribution of devices by country.
@@ 3.1.2 MonIoTr
To facilitate our cross-environment evaluation, we define four distinct data partitions from this dataset based on the physical location of the device and its network routing. This was made possible as the original data collection included experiments where traffic was routed between sites using a VPN. The four partitions are:
MonIoTr-UK: Data from devices physically located and operating on the UK network.
MonIoTr-USA: Data from devices physically located and operating on the USA network.
MonIoTr-UK-VPN: Data from devices physically located in the UK, but with their traffic routed to the USA site via a VPN before accessing the internet.
MonIoTr-USA-VPN: Data from devices physically located in the USA, but with their traffic routed to the UK site via a VPN.
@@ 3.2 Feature Extraction
4 Feature Selection
During the feature selection phase, we aimed to identify the most effective features for DI. For this, we used the four UNSW data partitions described in Section 3.1.1 to guide feature selection. These comprise two sessions (S1 and S2) from each of UNSW-AD and UNSW-DI (referred to hereafter according to their dataset-session number: AD-S1, AD-S2, DI-S1, and DI-S2).
The generalizability of features is assessed in the following three ways, in order of increasing strictness:
5-fold cross-validation (CV)
@@ 4 Feature Selection
4.1 Predictive Features
First, each feature is evaluated individually using each relevant combination of partitions, leading to 16 distinct evaluation contexts (see Figure 4). This process began with an initial pool of 332 features extracted as described in Section 3.2. Each of these 332 features was evaluated individually. For clarity, Figures 5 and 6 visualize only the subset of features (approximately 80) that demonstrated a non-zero kappa score in at least one of the 16 evaluation contexts. The remaining  250 features showed no predictive ability in any context and were thus eliminated from further consideration. From this reduced set of potentially useful features, we then applied our stricter voting mechanism to select the 46 most promising candidates for the GA-based interaction analysis. Generalizability is assessed by training a decision tree (DT) model using the feature, with DT selected for speed and explainability. This selection step, particularly when embedded within a GA, requires thousands of individual model training runs. Using a computationally inexpensive model like a DT was therefore essential to make the search process feasible. A more complex model (such as a random forest) would have been prohibitively time-consuming in this phase. The utility of each feature is measured using the kappa metric. The kappa statistic is a measure of inter-rater reliability that evaluates the agreement between a classifier and the ground truth, correcting for the probability that agreement could occur by chance.
@@ 4.1 Predictive Features
Figure 5 shows the resulting utility values. Notably, the scores measured under CV are much higher for some features than those measured under DD and SS scores. This indicates that information leaks may be occurring due to train and test folds being taken from the same partition, and suggests that CV is not a reliable basis for selecting features. Hence we only use DD and SS from now on.
@@
A voting system was used for feature selection, with each non-zero kappa value (with a tolerance of ±$\pm 5\%$, so $\geq 0.05$), indicating the feature had a positive contribution, resulting in a vote. Figure 6 illustrates this system. Features receiving four or more votes from either DD (green) or SS (red) categories, with at least one DD vote, advanced to the next stage. Features not meeting this criterion were eliminated. The requirement for at least one DD vote is based on its greater strictness in assessing generalizability. As a result of voting, we identified 46 of the 332 features as being individually predictive.
@@
Figure 4: (Left) Data usage from UNSW datasets across different evaluation contexts. CV, SS, and DD steps are visualized in shades of blue, red, and green, respectively, throughout the paper for clarity. In the nomenclature, CV|AD-S1 denotes cross-validation on the AD dataset session 1. For cases other than CV, the first dataset is used for training and the second for testing. For example, AD-S1|DI-S2 means the first session of the UNSW-AD dataset is used for training and the second session of the UNSW-DI dataset is used for testing. (Right) The legend displayed is shared across Figures 5-7 for consistency and clarity.
@@
Figure 5: Comparison of feature utility in UNSW datasets measured using CV (blue) and isolated methods, SS (red) and DD (green). CV tends to overestimate feature utility, with higher scores for many attributes. SS and DD produce more realistic evaluations. The discrepancy highlights the potential for information leakage in cross-validation and the importance of using isolated validation methods for assessing feature utility in ML-based DI models. For the legend, please refer to Figure 4.
@@
Figure 6: Features identified as being predictive within one or more evaluation contexts (CV: blue, SS: red, DD: green), based on non-zero kappa scores of DT models. Note that the CV votes are not used to select features, but are shown here for completeness. For the legend, please refer to Figure 4.
@@
Figure 7: List of intersecting features identified across eight dataset versus dataset (DD) cases using Genetic Algorithm (GA). Feature names highlighted in dark green were selected for model building after further post hoc analysis. For the legend, please refer to Figure 4.
4.2 Feature Interactions
We then considered feature interactions, using a wrapper method to find the optimal combination of these individually predictive features. We chose a Genetic Algorithm (GA) for this purpose, as its global search characteristic is effective at navigating large feature spaces to find robust solutions. A detailed description of the GA implementation, including the specific algorithm, hyperparameters, and evaluation criteria, is provided in Appendix Section G, Table 15.